Circular No. 7795

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001ij, 2001ik, 2001il, 2001im, 2001in A. V. Filippenko and R. Chornock, University of California, Berkeley,

report on inspection of CCD spectra (range 330–1000 nm), obtained on Jan. 17 UT with the Keck-I telescope, of the supernovae reported on IAUC 7778: “SN 2001ij is of type II-P at redshift 0.038, with hydrogen Balmer and Fe II lines having P-Cyg profiles; the object is still relatively bright. SN 2001ik is probably a supernova at redshift 0.094; some broad undulations appear to be present in its spectrum, which is heavily contaminated by light from the host galaxy. SN 2001il is of type Ia, roughly 2 months past maximum brightness, at redshift 0.208. SN 2001im, at redshift 0.075, may be of type IIb; He I lines appear to be present, along with Hα absorption, but the spectrum is too noisy for definitive classification. The putative host of SN 2001in is at redshift 0.218, but the supernova is projected quite far from it. Moreover, the spectrum of SN 2001in more closely resembles that of a type-Ia supernova at redshift ≈ 0.35. Accurate color measurements near the time of discovery should help resolve this discrepancy. The redshifts of the host galaxies are measured from narrow emission lines.”

XTE J1550–564 S. Corbel, Université Paris 7 and CEA, Saclay; R. Fender, University

of Amsterdam; and A. Tzioumis, Australia Telescope National Facility, report that Australia Telescope Compact Array radio observations of XTE J1550–564, obtained on Jan. 17, confirm the reappearance of radio emission associated with the recent reactivation in x-rays (IAUC 7792), with flux densities of 2.4 ± 0.1 and 2.5 ± 0.1 mJy at 4.8 and 8.6 GHz, respectively. They add: “This flat spectrum is as expected for a self-absorbed jet in the low/hard x-ray state. We see also an unexpected additional radio source, located at α = 15^{h}50^{m}55^{s}.97, δ = −56^{o}28′33′′.5 (equinox 2000.0; uncertainty < 1′′), > 20′′ west of the x-ray binary, with preliminary flux densities of 4.2 ± 0.1 mJy at 4.8 GHz and 2.4 ± 0.1 mJy at 8.6 GHz — consistent with an optically thin synchrotron spectrum. Systematic uncertainties in the flux densities of both sources may be larger than the formal errors, due to the proximity of a bright supernova remnant. The new source was not visible in images of the field in June 2000 or Feb. 2001, to a limit of 0.27 mJy (3σ) at 8.6 GHz, so it is transient; although the angular separation between it and XTE J1550–564 is large, we cannot rule out an association with the x-ray binary, although it may be an unrelated field source in coincidental outburst, in which case it may be affecting the x-ray data.”

2002 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green